0 out of 3 points Question 11 Sutton Company produces flash drives for computers, which it sells for $20 each. Each flash drive costs $6 of variable month. How much does Suttons operating income increase for each $1,000 increase in revenue per month? Selected Answer: c) $14,000 Answers: ·D costs to make. During April, 700 drives were sold. Fixed costs for April were $4 per unit for a total of $2.800 for the note that the University Bulet me and two hours of per week of the students a) $700 b) $50 c) $14,000 d) Not enough information to determine the answer DATE TOPICS Review of Intro to Organizat An Overview of d 04/02 04104 04109 The Changing Question 12 3 out of 3 points 0

need help please
Answer #1

Contribution margin ratio = 14*100/20 = 70%

Change in operating income = 1000*70% = 700

So answer is a) $700